SUB-DOPPLER SPECTROSCOPY OF THE Ã^{1}B_{1} - X̃^{1}A_{1} ELECTRONIC TRANSITION OF CBr_{2}

EYAD H AL-SAMRA and \underline{COLIN WESTERN}, School of Chemistry, University of Bristol, Cantocks Close, Bristol BS8 1TS, UK.

Sub-Doppler spectra of selected bands of the Ã^{1}B_{1} - X̃^{1}A_{1} transition of CBr_{2} are presented^{a} allowing all three rotational constants to be determined for the zero point level of the ground state for the first time. Refined rotational constants are presented for various Ã state vibrational levels, and an additional progression is identified in the Ã-X̃ absorption spectrum. The assignment of this progresion was assisted by modelling Franck-Condon factors with the pgopher^{b} program.

aE.H. Al-Samra and C.M. Western, J. Mol. Spectrosc. (2010), In Press, doi:10.1016/j.jms.2010.02.001

bPGOPHER, a Program for Simulating Rotational Structure, C. M. Western, University of Bristol, http://pgopher.chm.bris.ac.uk